# VAL-Bench: Measuring value alignment in Language Models

### v2

@@ 4.1 Metrics
We use Pairwise Alignment Consistency (PAC) as our main metric. For each input, we prompt the model twice: once with a “for” framing and once with an “against” framing. If the model answers consistently (e.g., supporting one side and opposing the other), it scores high. If the answers contradict each other or flip position, it scores low. This approach fits value alignment because there is often no single “correct” answer to compare against, but on the other hand, it’s possible for a model to score high while prioritizing values that many regard as undesirable. PAC works across many domains, avoids cherry-picking single outputs, and produces a clear score between 0 and 100 that can be calibrated across evaluators.
@@
Table 3 describes the notation for all the metrics we list in the results section. We do not define a preference for ideal NINF rate, since both overuse (becoming a substitute for refusals) and underuse (causing hallucinations) of no-information responses are problematic. Instead, we report it as an auxiliary metric, enabling future work to interpret its desirability in context.
Table 3: Definition and description of metric symbols used to form a comprehensive picture of alignment and its tradeoff with expressivity. Arrows denote the preferred direction of performance.
Symbol Definition Description
PAC $\uparrow$ $\frac{(\overline{{\mathbb{S}}}+2)\cdot 100}{4}$ Pairwise Alignment Consistency %
1REF $\downarrow$ $\tfrac{100}{N}\sum_{(r_{i}^{+},r_{i}^{-})\in{\mathbb{G}}_{\mathcal{M}}}\mathbf{1}\!\left[\psi(r_{i}^{+})+\psi(r_{i}^{-})=1\right]$ One side refusal rate %
2REF $\downarrow$ $\tfrac{100}{N}\sum_{(r_{i}^{+},r_{i}^{-})\in{\mathbb{G}}_{\mathcal{M}}}\mathbf{1}\!\left[\psi(r_{i}^{+})+\psi(r_{i}^{-})=2\right]$ Both side refusal rate %
1NINF $\tfrac{100}{N}\sum_{(r_{i}^{+},r_{i}^{-})\in{\mathbb{G}}_{\mathcal{M}}}\mathbf{1}\!\left[\nu(r_{i}^{+})+\nu(r_{i}^{-})=1\right]$ One side no-information response rate %
2NINF $\tfrac{100}{N}\sum_{(r_{i}^{+},r_{i}^{-})\in{\mathbb{G}}_{\mathcal{M}}}\mathbf{1}\!\left[\nu(r_{i}^{+})+\nu(r_{i}^{-})=2\right]$ Both side no-information response rate %
4.2 Evaluator calibration
@@
Because this evaluation only requires detecting consistency rather than value judgments, it is well-suited to automated scoring by an LLM. Despite the drastically increased objectivity of the task, it requires comprehension and analysis of complex opinions expressed in natural language. To ensure reliability, we create four calibration datasets by synthesizing responses constructed to target a certain behavior and measure the evaluator scores against human annotations on those datasets. Details of how we created these datasets can be found in the Appendix.
Table 4: Evaluator scores on calibration datasets. $\widehat{\cdot}$ denotes LLM-judge (Eval) scores; denotes human annotations on a subsample (Human). The calibration process provides confidence that the measurements are dependable.
Name Description $\widehat{\text{PAC}}$ $\widehat{\text{REF}}$ PAC REF
unaligned-0 Unaligned pairs, no refusals
Expected: PAC, REF = 0, 0
8.14 0.14 3.00 1.00
aligned-100 Aligned pairs, no refusals
Expected: PAC, REF = 100, 0
81.53 0.09 99.00 1.00
refusals-100 Both responses are refusals
Expected: PAC, REF = 100, 100
92.10 85.18 100.00 100.00
one-refusal-50 Exactly one of the two responses is a refusal
Expected: PAC, REF = 50, 50
46.49 43.32 51.00 51.00
Table 4 contains the description of the calibration datasets alongside measured PAC and REF scores by our chosen evaluator model and prompt (denoted by $\mathbf{\widehat{PAC}}$ and $\mathbf{\widehat{REF}}$), and by humans for whom we sampled 50 samples from each of the calibration datasets for a total of 200 human annotations (denoted by PAC* and REF*).
4.3 Evaluator ablations
We ran two ablations to ensure the accuracy of evaluation scores. Since the ordering of information in the input to LLMs matters, we conducted the evaluation with the paired responses presented in reverse order. To measure the deviation, we computed RMSE using both sets (normal and reversed orderings) of PAC, REF, and NINF metrics, resulting in values of 2.28, 0.30, and 0.57, respectively. We report the mean of these two runs for each model in the results.
In another ablation, we generated paraphrased versions of the responses in the unaligned-0 and aligned-100 calibration datasets (using an LLM instructed to make the responses concise and increase readability), and recomputed the metrics. This resulted in a PAC score increase of 3.71 and 3.63 for unaligned-0 and aligned-100 datasets, respectively. This indicates that the evaluator is relatively insensitive to changes in the style of expression.
Table 5: Chat model alignment scores sorted by alignment (metric definitions can be seen in Table 3). The alignment-expressivity tradeoff is clearly demonstrated in these results.
Arrows in the column headers denote the preferred direction of performance. Bold indicates best scores; underline indicates worst scores in each column.
Model PAC $\uparrow$ 1REF $\downarrow$ 2REF $\downarrow$ REF $\downarrow$ 1NINF 2NINF NINF
claude-haiku-3.5 68.79 30.38 15.14 30.34 16.41 9.55 17.76
claude-sonnet-4 68.34 8.70 2.37 6.71 17.01 15.02 23.52
claude-opus-4.1 66.11 9.57 3.22 8.01 14.24 9.11 16.23
claude-sonnet-3 55.27 13.97 1.89 8.87 14.04 2.46 9.49
qwen3-235B-instruct-2507 42.86 2.41 0.27 1.48 2.38 0.25 1.44
glm-4.5-air-nothink 34.38 5.14 0.32 2.88 11.18 1.29 6.88
llama-4-maverick-instruct 32.62 11.66 0.49 6.33 16.04 2.48 10.50
llama-3.3-70B-instruct 31.59 1.46 0.12 0.84 3.86 0.48 2.42
llama-2-70B-chat 30.43 6.61 0.16 3.46 0.15 0.00 0.08
gpt-3.5-turbo 26.91 0.89 0.04 0.48 0.98 0.09 0.57
llama-4-scout-instruct 26.43 12.44 0.97 7.19 11.78 0.49 6.38
qwen3-30B-instruct-2507 25.16 1.48 0.08 0.82 0.89 0.02 0.46
gpt-4o 24.61 0.18 0.02 0.11 0.72 0.10 0.46
gpt-4 23.76 0.77 0.10 0.48 5.04 0.12 2.65
gpt-4.1 22.77 0.09 0.01 0.05 0.25 0.03 0.16
mistral-large-instruct-2411 20.00 0.23 0.01 0.12 0.91 0.01 0.46
gpt-5-chat 19.49 0.11 0.01 0.06 0.15 0.00 0.08
gpt-4.1-nano 19.34 0.73 0.01 0.36 1.73 0.10 0.96
mistral-small-instruct-2506 16.66 0.30 0.00 0.15 0.40 0.00 0.20
gpt-4.1-websearch 10.70 0.57 0.01 0.29 0.41 0.04 0.24
glm-4.5-air-base 7.42 0.28 0.00 0.14 0.37 0.00 0.23
@@ 4.4 Thinking and Chat models
We evaluate a gamut of LLMs accessible through API or open weights. Ever since the launch of OpenAI’s o1 (OpenAI et al. (2024)), models trained with RLVR (DeepSeek-AI et al. (2025)) to output a detailed chain-of-thought (CoT) before the final response have shown incredible performance. We measure whether these models can utilize CoT to improve value alignment. We refer to these models as reasoning or thinking models, and the “regular” models as chat models.
@@ 5 Results
5.1 Chat models
As shown in Table 5, Claude models score the highest on alignment, but also have correspondingly higher refusal rates and no-information response rates (particularly for claude-haiku-3.5).
OpenAI’s models and most open-weights models typically score far lower on the alignment leaderboard. qwen3-235B-instruct-2507 achieves the highest PAC score among non-Claude models, while having relatively low refusal and no-information response rates.
We also include a pretrained only model glm-4.5-air-base in the list. This is the model we used for generating the unaligned-0 calibration dataset. As one might expect, it invariably provided the abductive reasoning expected by the prompts, leading to the lowest alignment score.
5.2 Thinking models
With the ability to spend a considerable number of tokens dedicated towards CoT, models may be expected to achieve higher alignment. The results (shown in Table 6) don’t show a consistent pattern, however. Among the models tested, only qwen3’s thinking models show a meaningful improvement. OpenAI’s o4-mini performs the worst of any model in the gpt-4 series.
Table 6: Thinking model alignment score sorted by alignment (metric definitions can be seen in Table 3). Qwen3 thinking models improve alignment compared to their non-thinking counterparts.
Arrows in the column headers denote the preferred direction of performance. Bold indicates best scores; underline indicates worst scores in each column.
Model PAC $\uparrow$ 1REF $\downarrow$ 2REF $\downarrow$ REF $\downarrow$ 1NINF 2NINF NINF
claude-sonnet-4-thinking 68.26 8.64 2.34 6.66 17.11 14.96 23.52
qwen3-235B-thinking-2507 57.42 2.74 0.27 1.63 2.91 0.32 1.77
qwen3-30B-thinking-2507 49.34 2.46 0.28 1.51 3.00 0.23 1.73
glm-4.5-air-thinking 33.59 1.60 0.12 0.93 14.86 4.63 12.06
gpt-5 26.41 0.85 0.03 0.46 0.72 0.06 0.42
deepseek-r1 12.61 0.17 0.01 0.10 0.17 0.00 0.09
o4-mini 12.04 6.30 0.08 3.24 0.10 0.00 0.04
@@ 5.3 Values demonstrated in responses
Using the value annotations of the prompts (Section 3.2), we directly analyze the values models demonstrate. Specifically, we use an LLM to annotate each response with the values it demonstrates from among those inspired by the prompt. Remember, this value expectation is implicit - the models demonstrate these values naturally.
To understand a model’s value system, we measure deviation in the frequency distribution of expected values (extracted from prompts) and demonstrated values, using the established value taxonomy. The expected distribution, derived from both for and against prompts, represents an inconsistent value system, and high deviation indicates a well-defined value system. We plot the Pearson residuals (Figure 2) for the ten highest residual values (denoting over-representation) and the ten least residual values (denoting under-representation).
Even with glm-4.5-air-base, a pretrained model with no post-training, the responses show a shift towards many morally good values (prioritizing accountability, prioritizing equity) and against many morally bad values (evading accountability, marginalization). But, the list of over-represented values includes some morally ambiguous ones, like pragmatism over principle and utilitarianism. gpt-5-chat’s ten highest residual values also include pragmatism over principle. claude-opus-4.1 seems to have the largest absolute residuals among the three, while having subtly different values in the top ten highest/least residuals.
@@
VDRVDR is the Value Demonstration Rate, measuring value expressivity. $\tilde{\chi}^{2}$ is the Reduced ${\chi}^{2}$ Statistic; indicating value alignment. These metrics illustrate the alignment-expressivity tradeoff emerging from direct value analysis.
To summarize the trend, we define value demonstration rate ($VDR$) as the ratio of the sum of frequencies of demonstrated values compared to the sum of frequencies of expected values, and use reduced $\chi^{2}$ statistic ($\tilde{\chi}^{2}$) as an aggregate measure of the deviation. Both are shown in Figure 2, revealing the alignment-expressivity tradeoff.
5.4 Correlations with issue awareness
We expect model alignment to differ depending on the kind of issue that’s being prompted. We measured how alignment rates (PAC), refusal rates (REF), and no-information response rates (NINF) vary with Issue awareness. The results are shown in Figure 3.
The correlations are quite clear: NINF shows a negative correlation, matching our intuition that as issue awareness increases, models avoid the no-information response. Claude models’ propensity to use no-information responses was also observed in the SimpleQA benchmark (Wei et al., 2024).
(a) No-info rate ($\overline{\rho}=-0.99$)
(b) Refusal Rate ($\overline{\rho}=0.96$)
(c) Alignment Score ($\overline{\rho}=0.89$)
Figure 3: Metrics vs Issue awareness. We see clear correlations with all 3 primary metrics.
PAC and REF show positive correlations. While refusal rates increase with issue awareness, the positive correlation with alignment itself exists independently of refusal rates. Both gpt-5-chat and deepseek-r1 show that trend while having negligible refusal rates. It suggests a bias toward over-prioritizing prominent issues in datasets used in alignment training of these models.
@@ 6 Discussion
Our VAL-Bench results highlight the dependence of alignment on the internal choices of each frontier model builder. OpenAI models consistently score low while avoiding refusals. Claude shows significantly higher alignment alonside higher refusal or no-information response rates. This dichotomy clearly illustrates the alignment-expressivity tradeoff, or alignment tax (Ouyang et al., 2022). Qwen3 (Yang et al., 2025) appears to pursue alignment through explicit opinions with low refusal rates. Given AI alignment’s societal importance and varying interpretations among frontier labs, we need a common, open process for creating human-annotated alignment datasets.
6.1 Training for value alignment
@@
All LLMs go through a large-scale pretraining phase with the cross-entropy objective that can be derived as the Maximum Likelihood Estimator. Modern LLMs are very effective at reducing this loss (Dubey et al., 2024), thereby modeling the broad distribution of viewpoints present in their training data, leading to unaligned models as illustrated by glm-4.5-air-base’s PAC score. It is impractical to “clean” the extraordinarily large pretraining datasets (up to 36T tokens, as per Yang et al. (2025)) to encode a consistent value system.
Alignment is commonly pursued through RL methods like RLHF (Ouyang et al. (2022)) or preference-based methods like DPO (Rafailov et al., 2023). Im & Li (2024) shows that the effectiveness of DPO depends heavily on large, diverse datasets, which helps explain why organizations invest in costly human feedback datasets (Lee et al. (2023), (Bai et al., 2022)). These methods require extensive datasets demonstrating correct behavior across many situations, and our results highlight the challenge of scaling value alignment while maintaining expressivity.
6.1.1 Belief consistency in language model training
In humans, beliefs often underpin values (Schwartz et al., 2012). Thus, achieving belief consistency also entails, at least in part, achieving value consistency. A training algorithm that is compatible with belief revision while maintaining consistency, like the AGM belief revision theory (Alchourrón et al., 1985), could, in principle, reduce the need for large and diverse datasets with explicit examples of alignment behavior; the consistent belief system would be formed during the self-supervised training phase itself. The variation and inconsistency of beliefs in pretraining datasets could actually inform the belief entrenchment (e.g., “vaccines are good for society” is strongly held, whereas “technology is an inherent good” is weakly held), as proposed in the AGM theory (Makinson & Gärdenfors, 2015). However, currently, we lack operators that measure logical relationships like contradiction or consequence in neural representation spaces (Garcez & Lamb, 2023), preventing direct implementation of such a method.
7 Conclusion
We introduced VAL-Bench, a large-scale benchmark testing whether LLMs demonstrate consistent values across prompts with opposing framings of controversial issues. Unlike prior benchmarks focused on predefined safety violations like toxicity, VAL-Bench directly measures value expression coherence. Leading open and closed-weights models show substantial variation in value alignment consistency; models relying heavily on refusals often avoid expressing values, while more expressive models risk inconsistency across framings. These trade-offs highlight a core challenge in alignment: safety mechanisms can mask, but not resolve, whether models embody stable value systems. By grounding evaluation in real-world controversies and enabling scalable automated assessment, VAL-Bench provides a reproducible tool for tracking value alignment progress and building models with consistent value systems. Future work could extend VAL-Bench beyond English Wikipedia, incorporate more universal value taxonomies, and investigate mechanistic causes of inconsistency.
Ethics Statements

### v3

@@ 4.1 Metrics
Table 3 describes the notation for all the metrics we list in the results section. We do not define a directional preference for ideal NINF rates as well, since both overuse (becoming a substitute for refusals) and underuse (causing hallucinations) of no-information responses are problematic in different ways. Instead, we report it as an auxiliary metric, enabling future work to interpret its desirability in context.
4.2 Evaluator accuracy
@@
Because this evaluation only requires semantic analysis rather than value judgments, it is well-suited to automated scoring by an LLM-as-a-judge. Despite the drastically increased objectivity of the tasks, it still requires comprehension and analysis of complex opinions expressed in natural language.
To compute evaluator accuracy, we create a human annotation task that requires them to annotate the same labels on the same scales (binary or Likert) as the LLM. The annotation prompt set was created by uniformly sampling response pairs from 5 language models developed by different model developers, yielding 200 samples. The annotation was performed independently by two of the authors. Table 4 summarizes the agreement stats, with Cohen’s $\kappa$ the main agreement measure (we use quadratic weighting for PAC).
The LLM achieves human-level agreement across most metrics, with human-LLM agreement scores matching or bettering human-human baselines. VPREF (Value Preference) is an exception, showing a noticeable bias by under-reporting value preference as compared to humans.
Table 4: Human-LLM evaluator agreement on calibration datasets. H represents Human, $\mathcal{E}$ represents LLM Evaluator
Metric N Rate-H (Avg) HH Agr. $H$ $\kappa$ Rate-$\mathcal{E}$ H-$\mathcal{E}$ Agr. (%) H-$\mathcal{E}$ $\kappa$
PAC 200 48.25 % - 0.68 47.75 % - 0.81, 0.70
VPREF 200 38.00 % 76 % 0.49 19.50 % 77 %, 70 % 0.46, 0.30
REF 400 13.00 % 87 % 0.43 13.75 % 95 %, 87 % 0.75, 0.49
NINF 400 9.38 % 97 % 0.78 8.5 % 98 %, 96 % 0.89, 0.76
We annotated a separate training set to optimize the evaluator. We use Gemma-3-27B-it as the LLM, use nucleus sampling with $temperature=0.05$ and top_p=0.9top\_p=0.9. The prompt is in the appendix.
@@ 4.3 Thinking and Chat models
We evaluate a gamut of LLMs accessible through API or open weights. Since the launch of OpenAI’s o1 (OpenAI et al. (2024)), models trained with RLVR (DeepSeek-AI et al. (2025)) to output many tokens of chain of thought before the final response have shown incredible performance on many complex benchmarks. It remains unclear how RLVR affects alignment, and we attempt to measure this impact by using their non-reasoning counterparts and comparing the results. We refer to these models as reasoning or thinking models, and the other models as chat models.
@@ 5 Results
Table 5: VAL-Bench scores sorted by PAC (metric definitions can be seen in Table 3). The consistency-expressivity tradeoff is seen in the correlation between PAC and REF/NINF ($\rho=0.91$).
Arrows denote the preferred direction of performance. Bold indicates best scores.
Model PAC $\uparrow$ VPREF REF $\downarrow$ NINF
Chat models
claude-haiku-3.5 81.51 27.63 47.84 13.13
claude-opus-4.1 79.10 25.14 38.70 15.54
claude-sonnet-4 79.00 25.64 30.68 23.73
claude-sonnet-3 64.93 29.93 18.54 7.40
llama-4-scout-instruct 51.09 22.99 31.24 1.47
llama-4-maverick-instruct 48.32 22.76 18.76 7.12
qwen3-235B-instruct-2507 46.77 26.63 5.77 1.24
gpt-4 45.05 4.85 27.00 0.56
glm-4.5-air-nothink 43.45 19.83 13.86 5.93
llama-3.3-70B-instruct 41.22 19.11 9.96 1.92
llama-2-70B-chat 34.77 22.64 11.71 0.03
gpt-3.5-turbo 31.27 12.19 4.00 0.40
gpt-4o 26.73 7.26 1.19 0.37
qwen3-30B-instruct-2507 26.69 16.14 2.68 0.22
gpt-4.1-nano 25.34 5.69 5.01 0.68
gpt-4.1 24.90 9.40 0.58 0.10
gpt-5-chat-latest 20.57 7.59 0.39 0.07
mistral-large-instruct-2411 23.08 8.11 2.40 0.40
mistral-small-3.2-instruct-2506 18.68 4.79 1.21 0.09
gpt-4.1-websearch 10.32 3.65 0.48 0.24
glm-4.5-air-base 10.26 2.41 2.30 0.03
Thinking models
claude-sonnet-4-thinking 78.90 25.43 30.50 23.63
qwen3-235B-thinking-2507 62.16 43.26 9.28 1.63
qwen3-30B-thinking-2507 53.25 36.76 7.10 1.62
glm-4.5-air-thinking 42.26 12.68 6.81 12.32
gpt-5 28.49 9.85 2.67 0.68
deepseek-r1 13.15 3.70 0.43 0.03
o4-mini 13.04 5.54 3.41 0.04
We summarize our main findings below, with full results in Tables 5.
Model to model variation. Models vary substantially in their scores ($\mu=41.4,\sigma=22.5$), suggesting that alignment training processes involve substantial subjective choices.
Pretrained model baseline. glm-4.5-air-base scores lowest on PAC, consistent with our intuition that pretrained models are not expected to exhibit belief consistency.
Only Claude models show high consistency. Belief consistency is the precondition for value alignment, and only Claude models get close to meeting that bar.
Driving factor for higher consistency. The tables show that both value preference and refusals (including no-information responses) correlate with consistency. The measured correlation of PAC with refusals (REF+NINF) is higher at $\rho=0.91$ than with value preference VPREF ($\rho=0.80$), indicating that consistency improvements are driven more by refusals than by value alignment.
Impact of CoT reasoning. Qwen3 thinking models show a significant increase in PAC compared to their instruct counterparts. In contrast to the overall trend, most of the increase is driven by value preference, making qwen3-235B-thinking-2507 the most value-aligned model among our results by far. The ability to use CoT to improve alignment is unique to Qwen.
@@ 5.1 Values demonstrated in responses
VDRVDR is the Value Demonstration Rate, measuring value expressivity. $\tilde{\chi}^{2}$ is the Reduced ${\chi}^{2}$ Statistic; indicating value preference. These metrics illustrate the alignment-expressivity tradeoff emerging from direct value analysis.
While VAL-Bench eschews actual value judgment, it’s possible to use the prompt and response dataset to describe the model’s value system. Using the value annotations of the prompts (Section 3.2), we directly analyze the values models demonstrate. Specifically, we use an LLM to annotate each response [^7] LLM annotations were not systematically verified. This analysis illustrates potential uses of this dataset beyond the primary evaluation metrics. with the values it demonstrates from among those encoded in the prompt. This value expectation is implicit - the models demonstrate these values naturally without needing to directly name the value in the prompt.
To provide a description of a model’s overall value system (value preferences that hold across contexts), we measure how the distribution of demonstrated values deviates from the distribution of expected values (extracted from prompts), using a common value taxonomy. The expected distribution is derived from both for and against prompts and thus, represents a neutral balance of values. Deviation from this expectation indicates value preference across contexts [^8] An agent showing no deviation is either very inconsistent, highly value indifferent, or a mix of the two.. We plot the Pearson residuals (Figure 2) for the ten highest residual values (denoting positive preference) and the ten least residual values (denoting negative preference).
Even glm-4.5-air-base, a pretrained model with no post-training, shows a preference for many morally good values (prioritizing accountability, prioritizing equity) and against many morally bad values (evading accountability, marginalization). But, the list of preferred values also includes some morally ambiguous ones, like pragmatism over principle and utilitarianism. gpt-5-chat’s ten highest residual values also include pragmatism over principle. claude-opus-4.1 seems to have the largest absolute residuals among the three with no clearly identifiable morally gray value in the top preferred values.
We also define value demonstration rate ($VDR$) as the ratio of the sum of frequencies of demonstrated values compared to the sum of frequencies of expected values, and use reduced $\chi^{2}$ statistic ($\tilde{\chi}^{2}$) as an aggregate measure of the deviation. Both are shown in Figure 2, and the anti-correlation between these two metrics reveals the alignment-expressivity tradeoff.
@@ 6 Discussion
It’s our hope that these results will motivate the community to consider belief consistency as an important goal of alignment training, especially for models powering general purpose chatbots. Lack of belief consistency manifests as expressions that agree with the viewpoint contained in the prompt. Moreover, modern LLMs are able to provide persuasive and personalized reasons justifying that viewpoint. It’s easy to see how this behavior might amplify biases of users or allow disinformation actors to scale their content strategy further. Chatbot app puts the burden of verification on the user [^9] Claude app uses the language “Claude can make mistakes. Please double-check responses.”, ChatGPT “ChatGPT can make mistakes.” and Gemini “Gemini can make mistakes, so double-check it.” Unfortunately, research consistently shows that users form beliefs from information they find online (Aslett et al. (2024), Allcott et al. (2024)), including AI chatbots, for which the effect is stronger due to the personalization of responses (Matz et al., 2024).
6.1 Training belief consistency
@@
All LLMs go through a large-scale pretraining phase with the MLE compatible objective to minimize cross-entropy loss for the next token. Modern LLMs are very effective at reducing this loss (Dubey et al., 2024), thereby modeling the broad distribution of viewpoints present in their training data. As the PAC score for glm-4.5-air-base shows, this process leads to inconsistent models. Any practical pretraining dataset will contain reasonable pluralism, and thus belief inconsistency in models is an expected outcome of pretraining.
Alignment is commonly pursued through RL methods like RLHF (Ouyang et al. (2022)) or preference-based methods like DPO (Rafailov et al., 2023). The current implementation of these methods, used by practically all models during post-training, doesn’t lead to consistency (as shown by our results). Preference datasets for contexts relating to divisive issues will likely contain a high level of disagreement. Research on the impact of such datasets with high disagreement is an active area of exploration (Zhang et al. (2024), Ali et al. (2025)), but none seem to have considered the goal of belief consistency yet.
Constitutional AI (CAI) (Bai et al., 2022) achieves value alignment by having the AI critique and revise its own outputs according to a set of written principles (the constitution), rather than relying solely on human labelers. A well drafted constitution that is applicable to a diverse set of issues can lead to belief consistency in a model.
7 Conclusion and Future Work
We introduced VAL-Bench, a large-scale benchmark for testing whether LLMs apply human values consistently across prompts with opposing framings of controversial issues. We make a concrete argument that belief consistency is a necessary requirement for value alignment. Empirically, we find that most LLMs exhibit a high amount of belief inconsistency and argue that due to the potential harms of such behavior, it should be prioritized as an important goal of alignment training. We also show that models with higher consistency achieve that primarily via refusals, lowering expressivity of those models in ambiguous and sensitive situations. By grounding evaluation in real-world controversies and enabling scalable, automated assessment, VAL-Bench provides a reproducible tool for tracking progress in value alignment training.
Future work could extend VAL-Bench into other languages. Languages form natural boundaries in communication, and thus they may foster a unique ecosystem of values. Applying these methods across different languages could be very instructive. Another natural extension is to evaluate word-deed consistency. LLMs are increasingly being provided more and more agency through the use of tools, and it is important for them to do what they say.